OpenStudy (anonymous):

Question about completing squares? What value completes the square x2 + 14x?

OpenStudy (anonymous):

Rediculous as it may sound, not sure I get this question..e_e

OpenStudy (anonymous):

14/2 = 7

OpenStudy (anonymous):

(x+7)(x+7) = x^2 +14x + 49